OpenStudy (ksaimouli):

double integral y dA bounded by y=x-2, x=y^2

OpenStudy (ksaimouli):

\[\int\limits_{}^{}\int\limits_{}^{} y dA\]

OpenStudy (ksaimouli):

in dxdy and dydx how to get the limits

ganeshie8 (ganeshie8):

sketch the region

OpenStudy (ksaimouli):



ganeshie8 (ganeshie8):

any reason you're not showing y = -sqrt(x) part ?

ganeshie8 (ganeshie8):